Circular No. 6946

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cl IN MARKARIAN 261 A. V. Filippenko, D. C. Leonard, and M. Modjaz, University of Cali-

fornia at Berkeley; and R. G. Eastman, Lawrence Livermore National Laboratory, report that CCD spectra (range 320–1000 nm) obtained on June 18 UT with the Shane 3-m reflector at Lick Observatory show that SN 1998cl (cf. IAUC 6934) is of type Ia, roughly 2 weeks past maximum brightness.

MICROLENSING EVENT N. Palanque-Delabrouille, Centre d’Etudes de Saclay, on behalf on

the EROS collaboration (cf. IAUC 6609), reports on observations of the second caustic crossing of the exotic gravitational microlensing event in progress toward the Small Magellanic Cloud (cf. IAUC 6935, 6939): “The 40 data points collected at the European Southern Observatory, La Silla, with the 1-m EROS telescope span the period June 18.23–18.45 UT. A linear decrease of the flux by about a factor of 5 occurred until June 18.31, followed by a slower rise by a factor of 2 in ∼ 3 hr. This is compatible with the predictions made by both the MACHO and the PLANET (cf.

http://www.astro.rug.nl/~planet/index.html) collaborations. Light

curves will be posted at our web site http://www.lal.in2p3.fr/EROS/ erosa.html as they become available.”

SUPERNOVA 1998bw IN ESO 184-G82 T. P. Dominici, R. Teixeira, and J. E. Horvath, Instituto Astronômico

e Geofísico (IAG), Universidade de São Paulo (USP); and P. Holvorcem, Instituto de Matematica e Estatistica, Universidade Estadual de Campiñas, report: “The possible optical counterpart (IAUC 6895, 6901, 6903) for GRB 980425 (IAUC 6884) has been observed at the IAG/USP Abrahão de Moraes Observatory (Valinhos, Brazil). Observations were performed with the 0.19-m Askania-Zeiss Meridian Circle equipped with a CCD camera on May 9.2, 10.2, 12.2, 19.2, 20.2, and 22.2 UT. Reduction of the whole data set using specific astrometric techniques (Teixeira et al. 1992, A.Ap. 264, 307) has provided the position α = 19^{h}35^{m}03^{s}.310 ± 0^{s}.007, δ = −52^{o}50′44′′.65 ± 0′′.05 (equinox J2000.0, FK5 reference frame). This refined position is remarkably close to that of the radio transient reported by Tinney et al. (IAUC 6896).”

1998 June 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green